# Shear stresses in fluid and solid membranes with bending elasticity

### v1

@@ Abstract
Comparison of a few simple models of fluid and solid membranes illustrates how shear stresses can arise from a bending energy through a coupling between curvature and surface stresses, a feature incidental to the fluid or solid nature of the material. In particular, it is shown how a fluid-like Helfrich bending energy contributes shear stresses, while a related solid-like energy, the correct continuum limit of a Seung-Nelson discrete bending term, produces a stress tensor with purely isotropic tangential part. An important distinction is noted between the resistance to tangential flows on the surface and the ability to support tangential stresses. The tangential projection of the divergence of the stress, or the pseudomomentum balance, is viewed in the light of some statements and observations in the literature. Different types of constraints, invariance properties, and shape equations are also briefly discussed.
@@ I Introduction
In this brief note, we consider bending energies extracted from two related membrane models commonly used in soft matter research, the Helfrich energy [1] based on squared mean curvature, and the continuum limit of the discrete Seung-Nelson (SN) energy [2] based on differences between normals on a triangular mesh. One is a fluid model that supports shear stresses, and the other is a solid model that does not— to be precise, the bending portion of the model does not support them, but it is usually accompanied by an additional stretching energy that does. For further illustration, we pair these bending energies with either a fluid-like constraint on area or a solid-like constraint on the metric. We observe that physicists often incorrectly identify the SN bending energy with Helfrich, an approximation that applies in the limit of small strain, only enforceable for sufficiently stiff or fully metrically-constrained solid membranes. As a result, the continuum form of this model has not been thoroughly examined. [^1] The name Seung-Nelson follows common convention in the soft matter field, but perhaps a more historically accurate name would be the Kantor-Nelson energy [3]. Kantor and Nelson make the reasonable identification of their continuum limit with a small-strain, small-slope approximate plate energy at lowest order in displacement derivatives, where the distinction between present and referential gradients and metrics is small. Seung and Nelson incorrectly infer a more general exact correspondence between the difference in normals and the (present) gradient of the normal field, and thereby with Helfrich. Seen within the context of their broad and insightful paper, the error is of minor and incidental significance, but the paper has been very influential; the downstream effects of this misidentification have at times been important, for example whenever the coupling between curvature and tension is of interest. An exhaustive list of soft matter papers where this error has propagated is impractical here. It makes a notable appearance in an early high-profile publication [4] (despite its title, this paper does not model a fluid membrane, but a bead-spring SN membrane), and again in a review by the same authors [5]. Many other works repeat the error, while many others do not explicitly make the error, but do not clearly address the approximation.
@@
That a fluid model can generate shear stresses has been a common point of confusion in the literature. While it is well known that the Helfrich energy produces an anisotropic tangential contribution to the membrane stress tensor, as derived for example in [6, 7, 8, 9, 10, 11], many users of this energy are uncomfortable identifying these terms as shear stresses, because of a misperception that such quantities cannot be supported by a fluid membrane (examples are provided below in Section II). There are two main points we wish to make. The first is that the quantities appearing are indeed shear stresses that can resist forces tangential to the membrane, but they arise from the resistance to bending, that is, deformations of the shape of the membrane, and not to shear deformations within the membrane that tangentially displace material. In a fluid film with bending elasticity, only the former deformations contribute to the energy. This identification of shear stresses is rarely stated explicitly (exceptions are provided below in Section II). The second point is that shear or any other tangential stresses are not some inevitable consequence of “geometry”, but need not arise at all from a bending energy. The appearance and sign of these terms are due to the dependence of the energy on the stretch, the tensor encoding the ratio of present to referential lengths, which dependence governs the coupling between bending and tangential stresses [12, 13, 14]. In particular, the fluid Helfrich bending energy displays this coupling, while the solid SN bending energy has a more primitive constitutive response that does not give rise to shear or any other tangential stresses (Section IV; what few new calculations are needed are presented in an appendix). The stretch is a quantity relevant in elasticity rather than fluid mechanics, and the insight just mentioned arises unexpectedly from considerations natural to the elasticity of thin solid bodies that may seem foreign to the study of fluid membranes. We round out the note with a minor point on the identification of projections of the momentum balance equations with the pseudomomentum balance (Section V), noting in passing how certain rearrangements of terms made possible by geometric relations serve to mask some of the structure and relationships present.
@@ II Fluidity and its malcontents
The energy of an incompressible fluid surface is invariant under area-preserving diffeomorphisms of some chosen reference configuration, or equivalently under the unimodular group of transformations of its tangent space [6, 7]. Physically, this corresponds to deformations consisting of area-preserving, shape-preserving (tangential) rearrangements of material on the surface. This fluidity is distinct from the ability to reparameterize (relabel) material elements in a homogeneous medium. The latter property is not unique to fluids, and gives rise to the conservation of pseudomomentum discussed in Section V. As shear deformations within a fluid surface do not cost energy, they produce no static forces, and will be resisted only by viscous forces if present. By contrast, rearrangements (but not reparameterizations) of a solid membrane are tightly constrained to preserve a two-dimensional metric, not just an area.
@@
Some researchers have recognized, at times reluctantly, the existence of shear (anisotropic) stresses. Powers and co-workers [23] begin by considering only the normal part of the momentum balance. Their assumption that a fluid membrane must have isotropic tension leads to an “apparent paradox”. Upon further comparison with force and moment balances for axisymmetric, nearly flat plates, they conclude that the bending moments contribute an additional “tension” (quotes theirs), and suggest a possible physical mechanism for its origins. Others derive tangential stresses from variation of the energy, in which anisotropic terms clearly appear. Tu and Ou-Yang [24] state that “we seem to arrive at a paradox… fluid membranes cannot withstand in-plane shear strain, however… shear stress still exhibits in non-spherical vesicles.” Guven and Vázquez-Montejo [25] state that “There will generally be a geometrical in-plane shear [force]… There is, of course, no inconsistency with the fluid character of the membrane.” Sauer and co-workers [26] state that “Interestingly, the bending part of the Helfrich model can contribute an in-plane shear stiffness”. Other, and some of the same, authors make incorrect statements regarding stresses, namely that the membrane offers no resistance to shear forces, or that what appear to be shear forces cannot actually be, because the membrane is fluid [5, 11, 18, 27, 28, 29]. Yet other, and some of the same, authors make ambiguous statements whose correctness depends on interpretation [30, 31, 26]. This representative sample of the works of prominent researchers in the field, including review articles and contributions to the 2016 CISM Advanced Summer School on “The Role of Mechanics in the Study of Lipid Bilayers”, should indicate the prevalence of the issue.
@@
We denote derivatives with respect to material coordinates as $d_{\alpha}$, $\alpha\in\{1,2\}$. These coordinates parameterize a surface ${\mathbf{x}}$ in $\mathbb{E}^{3}$ with normal ${\mathbf{n}}$, tangents ${\mathbf{a}}_{\alpha}=d_{\alpha}{\mathbf{x}}$, reciprocal tangents defined through the relations ${\mathbf{a}}^{\alpha}\cdot{\mathbf{a}}_{\beta}=\delta^{\alpha}_{\beta}$, metric components $a_{\alpha\beta}={\mathbf{a}}_{\alpha}\cdot{\mathbf{a}}_{\beta}$ and $a^{\alpha\beta}={\mathbf{a}}^{\alpha}\cdot{\mathbf{a}}^{\beta}$, area form $\textrm{d}a$, covariant derivative $\nabla_{\alpha}()$ constructed with the metric, and curvature components $b_{\alpha\beta}=b_{\beta\alpha}=d_{\beta}{\mathbf{a}}_{\alpha}\cdot{\mathbf{n}}=-{\mathbf{a}}_{\alpha}\cdot d_{\beta}{\mathbf{n}}$. Indices on these “present” quantities are raised and lowered with the metric. We denote the surface gradient $\nabla()=\nabla_{\alpha}(){\mathbf{a}}^{\alpha}$ and Laplacian $\nabla^{2}=\nabla^{\alpha}\nabla_{\alpha}$, and the mean and Gaußian curvature invariants $H=\tfrac{1}{2}b_{\alpha}^{\alpha}$ and $K=\tfrac{1}{2}\left(b_{\alpha}^{\alpha}b_{\beta}^{\beta}-b^{\alpha}_{\beta}b_{\alpha}^{\beta}\right)$.
@@
The parentheses are a reminder that the subscripts are not indices. Because of the constraints, these expressions represent the same energy, so the resulting shapes and stresses will be the same, but this requires different scalar constraints $\sigma_{\text{\tiny(H)}}$ and $\sigma_{\text{\tiny(rHa)}}$. Outside of the soft matter community, the bending part of (1) is often associated with the name of Willmore [32]. The full integral of this bending content— not its local density— is conformally invariant up to a boundary term [33], although this is not of much relevance in the presence of the area constraint. Besides inclusion of a referential curvature, Helfrich can also have different moduli for $H$ and $K$. The choice made here allows for easy comparison with SN and other elastic models, and is irrelevant to the points made in this paper. It is also common to consider a global constraint on area. In computations, it is more convenient to impose either global area constraints or a penalty term to avoid mixed finite element formulations, which are often sensitive to numerical instabilities. We will not delve into these options here.
@@
One should contrast $\nabla^{\alpha}{\mathbf{n}}\cdot\nabla_{\alpha}{\mathbf{n}}=\nabla^{\alpha}{\mathbf{n}}\cdot d_{\alpha}{\mathbf{n}}=a^{\alpha\beta}d_{\alpha}{\mathbf{n}}\cdot d_{\beta}{\mathbf{n}}=a^{\alpha\beta}a^{\gamma\eta}b_{\alpha\eta}b_{\gamma\beta}=b^{\alpha}_{\gamma}b^{\gamma}_{\alpha}$ from Helfrich with ¯$\bar{\nabla}^{\alpha}{\mathbf{n}}\cdot\bar{\nabla}_{\alpha}{\mathbf{n}}=\bar{\nabla}^{\alpha}{\mathbf{n}}\cdot d_{\alpha}{\mathbf{n}}=A^{\alpha\beta}d_{\alpha}{\mathbf{n}}\cdot d_{\beta}{\mathbf{n}}=A^{\alpha\beta}b_{\alpha}^{\gamma}b_{\gamma\beta}=A^{\alpha\beta}a^{\gamma\eta}b_{\alpha\eta}b_{\gamma\beta}=A^{\alpha\beta}a_{\gamma\eta}b_{\alpha}^{\eta}b^{\gamma}_{\beta}$ from SN (being careful to remember that indices contracted between $A^{\alpha\beta}$ and the other objects cannot be simultaneously raised and lowered). A physicists’ interpretation, namely the use of present or referential metric to raise indices on derivatives of the normal, can also be expressed in continuum mechanical language. Since the covariant components of the metric go like the square of stretch, comparison of the last terms shows that the SN energy goes like the square of the product of stretch and curvature. [^2] Continuum mechanicians might express the SN bending energy as the trace of either the present tensor ${\mathbf{b}}\cdot{\mathbf{B}}\cdot{\mathbf{b}}$ or, most likely, the referential tensor ${\bf{F}}^{\top}\cdot{\mathbf{b}}^{2}\cdot{\bf{F}}$, where ${\mathbf{b}}$ is curvature, ${\mathbf{B}}$ is the left Cauchy-Green deformation, and ${\bf{F}}$ is the deformation gradient, standard quantities that we do not define here. Similar re-expressions of quantities appearing in the Helfrich energy can be found in [7]. This difference is what governs the coupling of bending and in-plane stresses [12, 13, 14].
@@
Because of the constraints, these expressions represent the same energy, so the resulting shapes and stresses will be the same, but this requires different tensor constraints $\sigma_{\text{\tiny(rHm)}}^{\alpha\beta}$ and $\sigma_{\text{\tiny(SNm)}}^{\alpha\beta}$. Similarly, they are equivalent to the Guven-Müller functional [39]. In the context of an environment where one might find a lipid membrane, they might represent an unstretchable capsule.
@@
The stress (9) arising from the area-constrained Atluri/SN energy is tangentially isotropic. Aside from the normal term, its contraction with a tangent normal at the boundary (for example, $\nu_{\alpha}{\mathbf{f}}^{\alpha}$ where $\nu_{\alpha}$ is a component of the present tangent normal $\nu_{\alpha}{\mathbf{a}}^{\alpha}$) provides a purely tangent-normal (pressure-like) term and thus offers no resistance, whether through in-plane or bending deformations, to application of tangential shear forces. In contrast to the tensor multipliers in the two metric-constrained models (10-11), the scalar multiplier in the three area-constrained models (7-9) cannot absorb a tangentially anisotropic $-Hb^{\alpha\beta}$ term that appears in the Helfrich-type forces (7), (8), and (10) and provides a tangent-tangent contribution at the boundaries. This is the difference between constraining the area (9) and constraining the entire metric (11). These two models differ, as only the latter can provide constraint shear stresses.
The simple appearance in component form of the Atluri/SN stress (9) is deceptive. This is a hybrid quantity, of surprising complexity given the ease of computation of the discrete SN energy. As written, one applies the present covariant derivative (Christoffel symbols constructed with the present metric) to a quantity that “contracts” (yet the contracted indices cannot be simultaneously raised and lowered) the contravariant components of the referential metric (indices raised with the referential metric) with the mixed components of the curvature (index raised with the present metric). Some of this complexity is hinted at in Appendix A, where appear expressions involving a referential covariant derivative acting on a present tangent vector. [^3] Wrapping these quantities in index-free tensorial notation requires objects such as the two-point rotation or shifter tensors [40]. Physicists might be comfortable relating the latter to dot products between referential and present tangent vectors, a process some continuum mechanicians would reject as a mathematical abomination. In fact, the Atluri/SN energy is related to other quantities with similar stretch and curvature dependence that arise in elastic theories and are not easily expressible in terms of metrics and curvatures [36, 37].
@@
The tangential part of (12) is the gradient of a scalar or, equivalently, the divergence of an isotropic tensor. It is tempting to ascribe physical meaning to this, given that the Helfrich energy is fluid. This tensor also seems at first glance to arise only from the constraint. That the divergence of the bending contribution seems to not have a tangential component was noted in [9] and [42] (see also the appendix of [23], and [39]), with the latter invoking certain invariance properties of the energy as explanation, but this seems to be a conflation of reparameterization invariance and stationarity under tangential displacements of the sort criticized in [43]. Upon closer inspection, both of these features of the tangential term— isotropy and absence of bending contribution— may be physically irrelevant accidents for this particular geometric energy. It is potentially confusing when geometric quantities describe both the energy, and the geometry, of objects. The common arrangement of the term into the form (12), aided both by the Codazzi relations and the ability to move terms around that are proportional to pieces of the energy density, belies a more fundamental structure. There is an important distinction between the tangential part of the stress ${\mathbf{f}}^{\alpha}$ and the tangential part of the divergence of the stress. The tangential part of the stress will balance forces applied on boundaries, such as cuts in, or edges of, the membrane. The divergence of the stress appears in the balance of momentum, which has a tangential part; its tangential projection can be rearranged into the balance of pseudomomentum. In the current setting, this is a conservation law; the vanishing of the divergence of a “material stress” (also known as a pseudo-, Eshelby, or configurational stress, or the non-temporal part of an “elastic energy-momentum tensor” [44, 45, 46, 47, 48, 49, 50, 41]), is a consequence of the relabeling symmetry of a homogeneous material. External forces, whether body forces or forces applied from one or other side in the “interior” of the membrane, appear as source terms in the momentum balance but can be absorbed into the divergence in the pseudomomentum balance, whose source terms only come from broken material symmetries such as inhomogeneities in properties [51, 48, 50, 41].
@@
using an identity analogous to one mentioned earlier for Helfrich. This more clearly reveals the Eshelbian (elastic energy-momentum) form, in which the energy density appears subtracted from another quantity. This structure is always present, but here, unlike in the previous example, we cannot use Codazzi to simplify or rewrite in terms of gradients of scalar quantities. For completeness, the normal projection obtained from this model is

### v2

@@ Abstract
Comparison of a few simple models of fluid and solid membranes illustrates how shear stresses can arise from a bending energy through a coupling between curvature and surface stresses, a feature incidental to the fluid or solid nature of the material. In particular, it is shown how a fluid-like Helfrich bending energy contributes shear stresses, while a related solid-like energy, the correct continuum limit of a Seung-Nelson discrete bending term, produces a stress tensor with purely isotropic tangential part. A distinction is noted between the resistance to tangential flows on the surface and the ability to support tangential stresses. The tangential projection of the divergence of the stress, or the pseudomomentum balance, is viewed in the light of some statements and observations in the literature. Different types of constraints, invariance properties, and shape equations are briefly discussed. The state of a patch of unidirectionally-bent material is connected with prior analyses of the pulling of membrane tethers.
@@ I Introduction
In this paper, we consider bending energies extracted from two related membrane models commonly used in soft matter research, the Helfrich energy [1] based on squared mean curvature, and the continuum limit of the discrete Seung-Nelson (SN) energy [2] based on differences between normals on a triangular mesh. One is a fluid model that supports shear stresses, and the other is a solid model that does not— to be precise, the bending portion of the model does not support them, but it is usually accompanied by an additional stretching energy that does. For further illustration, we pair these bending energies with either a fluid-like constraint on area or a solid-like constraint on the metric. Physicists often incorrectly identify the SN bending energy with Helfrich, an approximation that applies in the limit of small strain, only enforceable for sufficiently stiff or fully metrically-constrained solid membranes. As a result, the continuum form of this model has not been thoroughly examined. [^1] The name Seung-Nelson follows common convention in the soft matter field, but perhaps a more historically accurate name would be the Kantor-Nelson energy [3]. Kantor and Nelson make the reasonable identification of their continuum limit with a small-strain, small-slope approximate plate energy at lowest order in displacement derivatives, where the distinction between present and referential gradients and metrics is small. Seung and Nelson incorrectly infer a more general exact correspondence between the difference in normals and the (present) gradient of the normal field, and thereby with Helfrich. Seen within the context of their broad and insightful paper, the error is of minor and incidental significance, but the paper has been very influential; the downstream effects of this misidentification have at times been important, for example whenever the coupling between curvature and tension is of interest. An exhaustive list of soft matter papers where this error has propagated is impractical here. It makes a notable appearance in an early high-profile publication [4] (despite its title, this paper does not model a fluid membrane, but a bead-spring SN membrane), and again in a review by the same authors [5]. Many other works repeat the error, while many others do not explicitly make the error, but do not clearly address the approximation.
@@
That a fluid model can generate shear stresses has been a common point of confusion in the literature. While it is well known that the Helfrich energy produces an anisotropic tangential contribution to the membrane stress tensor, as derived for example in [6, 7, 8, 9, 10, 11], many users of this energy are uncomfortable identifying these terms as shear stresses, because of a misperception that such quantities cannot be supported by a fluid membrane (examples are provided below in Section II). There are two main points we wish to make. The first is that the quantities appearing are indeed shear stresses that can resist forces tangential to the membrane, but they arise from the resistance to bending, that is, deformations of the shape of the membrane, and not to shear deformations within the membrane that tangentially displace material. In a fluid film with bending elasticity, only the former deformations contribute to the energy. This identification of shear stresses is rarely stated explicitly (exceptions are provided below in Section II). The second point is that shear or any other tangential stresses are not some inevitable consequence of “geometry”, but need not arise at all from a bending energy. The appearance and sign of these terms are due to the dependence of the energy on the stretch, the tensor encoding the ratio of present to referential lengths, which dependence governs the coupling between bending and tangential stresses [12, 13, 14]. In particular, the fluid Helfrich bending energy displays this coupling, while the solid SN bending energy has a more primitive constitutive response that does not give rise to shear or any other tangential stresses (this is shown in Section IV; what few new calculations are needed are presented in Appendix A, and a comparison for a unidirectionally-bent patch of material, as in certain models of pulled membrane tethers, is presented in Section VI). The stretch is a quantity relevant in elasticity rather than fluid mechanics, and the insight just mentioned arises unexpectedly from considerations natural to the elasticity of thin solid bodies that may seem foreign to the study of fluid membranes. Section V explores the identification of projections of the momentum balance equations with the pseudomomentum balance, noting in passing how certain rearrangements of terms, made possible by geometric relations, mask some of the structure and relationships present.
@@ II Fluidity and its malcontents
The energy of an incompressible fluid surface is invariant under area-preserving diffeomorphisms of some chosen reference configuration, or equivalently under the unimodular group of transformations of its tangent space [6, 7]. Physically, this corresponds to deformations consisting of area-preserving, shape-preserving (tangential) rearrangements of material on the surface. This fluidity is distinct from the ability to reparameterize (relabel) material elements in a homogeneous medium. The latter property is not unique to fluids, and gives rise to the conservation of pseudomomentum discussed in Section V. As shear deformations within a fluid surface do not cost energy, they produce no static forces, and will be resisted only by viscous forces if present. By contrast, rearrangements of a solid membrane are tightly constrained to preserve a two-dimensional metric, not just an area.
@@
Some researchers have recognized, at times reluctantly, the existence of shear (anisotropic) stresses. Powers and co-workers [23] begin by considering only the normal part of the momentum balance. Their assumption that a fluid membrane must have isotropic tension leads to an “apparent paradox” in the modeling of a membrane tether. Upon further comparison with force and moment balances for axisymmetric, nearly flat plates, they conclude that the bending moments contribute an additional “tension” (quotes theirs), and suggest a possible physical mechanism for its origins. Others derive tangential stresses from variation of the energy, in which anisotropic terms clearly appear. Tu and Ou-Yang [24] state that “we seem to arrive at a paradox… fluid membranes cannot withstand in-plane shear strain, however… shear stress still exhibits in non-spherical vesicles.” Guven and Vázquez-Montejo [25] state that “There will generally be a geometrical in-plane shear [force]… There is, of course, no inconsistency with the fluid character of the membrane.” Sauer and co-workers [26] state that “Interestingly, the bending part of the Helfrich model can contribute an in-plane shear stiffness”. Other, and some of the same, authors make incorrect statements regarding stresses, namely that the membrane offers no resistance to shear forces, or that what appear to be shear forces cannot actually be, because the membrane is fluid [5, 11, 18, 27, 28, 29]. Yet other, and some of the same, authors make ambiguous statements whose correctness depends on interpretation [30, 31, 26]. This representative sample of the works of prominent researchers in the field, including review articles and contributions to the 2016 CISM Advanced Summer School on “The Role of Mechanics in the Study of Lipid Bilayers”, should indicate the prevalence of the issue.
@@
We denote derivatives with respect to material coordinates as $d_{\alpha}$, $\alpha\in\{1,2\}$. These coordinates parameterize a surface ${\mathbf{x}}$ in $\mathbb{E}^{3}$ with normal ${\mathbf{n}}$, tangents ${\mathbf{a}}_{\alpha}=d_{\alpha}{\mathbf{x}}$, reciprocal tangents defined through the relations ${\mathbf{a}}^{\alpha}\cdot{\mathbf{a}}_{\beta}=\delta^{\alpha}_{\beta}$, metric components $a_{\alpha\beta}={\mathbf{a}}_{\alpha}\cdot{\mathbf{a}}_{\beta}$ and $a^{\alpha\beta}={\mathbf{a}}^{\alpha}\cdot{\mathbf{a}}^{\beta}$, area form $\textrm{d}a$, covariant derivative $\nabla_{\alpha}()$ constructed with the metric, and curvature components $b_{\alpha\beta}=b_{\beta\alpha}=d_{\beta}{\mathbf{a}}_{\alpha}\cdot{\mathbf{n}}=-{\mathbf{a}}_{\alpha}\cdot d_{\beta}{\mathbf{n}}$. Indices on these “present” quantities are raised and lowered with the metric. We denote the surface gradient and Laplacian as $\nabla()=\nabla_{\alpha}(){\mathbf{a}}^{\alpha}$ and $\nabla^{2}=\nabla^{\alpha}\nabla_{\alpha}$, and the mean and Gaußian curvature invariants as $H=\tfrac{1}{2}b_{\alpha}^{\alpha}$ and $K=\tfrac{1}{2}\left(b_{\alpha}^{\alpha}b_{\beta}^{\beta}-b^{\alpha}_{\beta}b_{\alpha}^{\beta}\right)$.
@@
Subscripts in parentheses are not indices. Because of the constraints, these expressions represent the same energy, so the resulting shapes and stresses will be the same, but this requires different scalar multipliers $\sigma_{\text{\tiny(H)}}$ and $\sigma_{\text{\tiny(rHa)}}$. Outside of the soft matter community, the bending part of (1) is often associated with the name of Willmore [32]. The full integral of this bending content— not its local density— is conformally invariant up to a boundary term [33], although this is not of much relevance in the presence of the area constraint. Besides inclusion of a referential curvature, Helfrich can also have different moduli for $H$ and $K$. The choice made here allows for easy comparison with SN and other elastic models, and is irrelevant to the points made in this paper. It is also common to consider a global constraint on area. In computations, it is more convenient to impose either global area constraints or a penalty term to avoid mixed finite element formulations, which are often sensitive to numerical instabilities. We will not delve into these options here.
@@
One should contrast $\nabla^{\alpha}{\mathbf{n}}\cdot\nabla_{\alpha}{\mathbf{n}}=\nabla^{\alpha}{\mathbf{n}}\cdot d_{\alpha}{\mathbf{n}}=a^{\alpha\beta}d_{\alpha}{\mathbf{n}}\cdot d_{\beta}{\mathbf{n}}=a^{\alpha\beta}a^{\gamma\eta}b_{\alpha\eta}b_{\gamma\beta}=b^{\alpha}_{\gamma}b^{\gamma}_{\alpha}$ from Helfrich with ¯$\bar{\nabla}^{\alpha}{\mathbf{n}}\cdot\bar{\nabla}_{\alpha}{\mathbf{n}}=\bar{\nabla}^{\alpha}{\mathbf{n}}\cdot d_{\alpha}{\mathbf{n}}=A^{\alpha\beta}d_{\alpha}{\mathbf{n}}\cdot d_{\beta}{\mathbf{n}}=A^{\alpha\beta}b_{\alpha}^{\gamma}b_{\gamma\beta}=A^{\alpha\beta}a^{\gamma\eta}b_{\alpha\eta}b_{\gamma\beta}=A^{\alpha\beta}a_{\gamma\eta}b_{\alpha}^{\eta}b^{\gamma}_{\beta}$ from SN (being careful to remember that indices contracted between $A^{\alpha\beta}$ and the other objects cannot be simultaneously raised and lowered). A physicists’ interpretation, namely the use of present or referential metric to raise indices on derivatives of the normal, can also be expressed in continuum mechanical language. As the covariant components of the metric go like the square of stretch, comparison of the last terms shows that the SN energy goes like the square of the product of stretch and curvature, [^2] Continuum mechanicians might express the SN bending energy as the trace of either the present tensor ${\mathbf{b}}\cdot{\mathbf{B}}\cdot{\mathbf{b}}$ or, most likely, the referential tensor ${\bf{F}}^{\top}\cdot{\mathbf{b}}^{2}\cdot{\bf{F}}$, where ${\mathbf{b}}$ is curvature, ${\mathbf{B}}$ is the left Cauchy-Green deformation, and ${\bf{F}}$ is the deformation gradient, standard quantities that we do not define here. Similar re-expressions of quantities appearing in the Helfrich energy can be found in [7]. rather than the square of curvature. The stretch dependence of the energy governs the coupling between bending and tangential stresses [12, 13, 14].
@@
Because of the constraints, these expressions represent the same energy, so the resulting shapes and stresses will be the same, but this requires different tensor multipliers $\sigma_{\text{\tiny(rHm)}}^{\alpha\beta}$ and $\sigma_{\text{\tiny(SNm)}}^{\alpha\beta}$. Similarly, they are equivalent to the Guven-Müller functional [39]. In the context of an environment where one might find a lipid membrane, they might represent an unstretchable capsule.
@@
The stress (9) arising from the area-constrained Atluri/SN energy is tangentially isotropic. Aside from the normal term, its contraction with a tangent normal at the boundary (for example, $\nu_{\alpha}{\mathbf{f}}^{\alpha}$ where $\nu_{\alpha}$ is a component of the present tangent normal $\nu_{\alpha}{\mathbf{a}}^{\alpha}$) provides a purely tangent normal (pressure-like) term and thus offers no resistance, whether through in-plane or bending deformations, to application of tangent tangent (shear) forces. In contrast to the tensor multipliers in the two metric-constrained models (10-11), the scalar multiplier in the three area-constrained models (7-9) cannot absorb a tangentially anisotropic $-Hb^{\alpha\beta}$ term that appears in the Helfrich-type forces (7), (8), and (10) and provides a tangent tangent contribution at the boundaries. This is the difference between constraining the area (9) and constraining the entire metric (11). These two models differ, as only the latter can provide constraint shear stresses. Section VI compares (7) and (9) for a rudimentary example, making use of further results from the following Section V.
The simple appearance in component form of the Atluri/SN stress (9) is deceptive. This is a hybrid quantity, of surprising complexity given the ease of computation of the discrete SN energy. As written, one applies the present covariant derivative (Christoffel symbols constructed with the present metric) to a quantity that “contracts” (yet the contracted indices cannot be simultaneously raised and lowered) the contravariant components of the referential metric (indices raised with the referential metric) with the mixed components of the curvature (index raised with the present metric). Some of this complexity is hinted at in Appendix A, where appear expressions involving a referential covariant derivative acting on a present tangent vector. [^3] Wrapping these quantities in index-free tensorial notation requires objects such as the two-point rotation or shifter tensors [40]. Physicists might be comfortable relating the latter to dot products between referential and present tangent vectors, a process some continuum mechanicians would reject as a mathematical abomination. In fact, the Atluri/SN energy is related to other quantities with similar stretch and curvature dependence that arise in elastic theories but are not easily expressible in terms of metrics and curvatures [36, 37].
@@
The tangential part of (12) is the gradient of a scalar or, equivalently, the divergence of an isotropic tensor. It is tempting to ascribe physical meaning to this, given that the Helfrich energy is fluid. This tensor also seems at first glance to arise only from the constraint. That the divergence of the bending contribution seems to not have a tangential component was noted in [9] and [42] (see also the appendix of [23], and [39]), with the latter invoking certain invariance properties of the energy as explanation, but this seems to be a conflation of reparameterization invariance and stationarity under tangential displacements, of the sort criticized in [43]. Upon closer inspection, both of these features of the tangential term— isotropy and absence of bending contribution— may be physically irrelevant accidents that occur with this particular geometric energy. It is potentially confusing when geometric quantities describe both the energy, and the geometry, of objects. The common arrangement of the term into the form (12), aided both by the Codazzi relations and the ability to move terms around that are proportional to pieces of the energy density, belies a more fundamental structure. There is an important distinction between the tangential part of the stress ${\mathbf{f}}^{\alpha}$ and the tangential part of the divergence of the stress. The tangential part of the stress will balance forces applied on boundaries, such as cuts in, or edges of, the membrane. The divergence of the stress appears in the balance of momentum, whose tangential projection can be rearranged into the balance of pseudomomentum. In the current setting, this is a conservation law; the vanishing of the (referential) divergence of a “material stress” (also known as a pseudo-, Eshelby, or configurational stress, or the non-temporal part of an “elastic energy-momentum tensor” [44, 45, 46, 47, 48, 49, 50, 41]), is a consequence of the relabeling symmetry of a homogeneous material. External forces, whether body forces or forces applied from one or the other side in the “interior” of the membrane, appear as source terms in the momentum balance but can be absorbed into the divergence in the pseudomomentum balance, whose source terms only come from broken material symmetries such as inhomogeneities in properties [51, 48, 50, 41].
@@
using an identity analogous to one mentioned earlier for Helfrich. This more clearly reveals the Eshelbian (material energy-momentum) form, in which the energy density appears subtracted from another quantity. This structure is always present, but here, unlike in the previous example, we cannot use Codazzi to simplify or rewrite in terms of gradients of scalar quantities. For completeness, the normal projection obtained from this model is